# Towards Scalable Exact Machine Unlearning Using Parameter-Efficient Fine-Tuning

### v1

@@ Abstract
Machine unlearning is the process of efficiently removing the influence of a training data instance from a trained machine learning model without retraining it from scratch. A popular subclass of unlearning approaches is exact machine unlearning, which focuses on techniques that explicitly guarantee the removal of the influence of a data instance from a model. Exact unlearning approaches use a machine learning model in which individual components are trained on disjoint subsets of the data. During deletion, exact unlearning approaches only retrain the affected components rather than the entire model. While existing approaches reduce retraining costs, it can still be expensive for an organization to retrain a model component as it requires halting a system in production, which leads to service failure and adversely impacts customers. To address these challenges, we introduce an exact unlearning framework – Sequence-aware Sharded Sliced Training (S3T), which is designed to enhance the deletion capabilities of an exact unlearning system while minimizing the impact on model’s performance. At the core of S3T, we utilize a lightweight parameter-efficient fine-tuning approach that enables parameter isolation by sequentially training layers with disjoint data slices. This enables efficient unlearning by simply deactivating the layers affected by data deletion. Furthermore, to reduce the retraining cost and improve model performance, we train the model on multiple data sequences, which allows S3T to handle an increased number of deletion requests. Both theoretically and empirically, we demonstrate that S3T attains superior deletion capabilities and enhanced performance compared to baselines across a wide range of settings. [^1] Work in Progress.
@@ 1 Introduction
In recent years, the growing success of machine learning (ML) has led to its widespread deployment across a range of applications [1, 58, 49]. Once a machine learning model has been trained, it may be necessary to unlearn specific training data instances for various reasons, like complying with user data deletion requests [43, 18, 52, 2], removing stale or corrupt data [5, 53], etc. Retraining an ML model entirely from scratch with each deletion request is expensive, especially for modern large-scale models [9, 1, 58]. Machine unlearning [46, 65] techniques focus on efficiently unlearning the influence of a data instance from a trained ML model.
@@
Machine unlearning techniques can be classified into two broad categories: approximate and exact unlearning. Approximate unlearning techniques [26, 40] modify the parameters of a trained model to reduce the influence of the deleted data instance. While cost-effective in practice, approximate unlearning cannot guarantee the complete removal of an instance’s influence and it may still retain non-zero influence on the model. Moreover, auditing approximate unlearning is challenging due to the stochastic nature of ML optimization [59]. An alternative to this approach is exact unlearning [10, 8], which can guarantee the removal of the influence of a data instance from a trained model. Exact unlearning techniques use a modular system, where different components within the system are trained on disjoint subsets of the data. When a deletion request occurs, only the affected component needs to be retrained. However, in real-world settings, halting a production system to even retrain a single component can result in service failure and significant expenses for a business. The alternative is to function without the affected component, which may result in reduced performance, ultimately impacting end users. To address these challenges, we introduce an efficient exact unlearning framework, Sequence-aware Sharded Sliced Training (S3T), which enhances the system’s deletion capability while minimizing the impact on performance.
@@
The key idea behind our S3T framework is to perform additional offline training before deploying the initial model to reduce retraining costs. At the core of S3T, we leverage a novel lightweight parameter-efficient fine-tuning (PEFT) that allows parameter isolation by sequentially training model layers using disjoint data slices. Due to this parameter isolation, it is possible to efficiently perform exact unlearning by deactivating the layers associated with the deleted instance, rather than discarding the entire checkpoint. We efficiently train multiple models using different sequences of the same data slices. We show that increasing the training budget before deployment can significantly reduce retraining costs and improve the model’s performance in production. We also observe that it is important to train the model using diverse sequences and provide an optimal algorithm for selecting such sequences using graph matching [13]. Furthermore, we formalize the notion of deletion rate for exact unlearning systems and theoretically show that S3T achieves provably better deletion guarantees than existing approaches. We conduct extensive empirical evaluations to evaluate the effectiveness of our proposed fine-tuning approach using Transformers with parameter counts ranging from 86M to 13B on a range of tasks. Additionally, we empirically validate the sequence selection algorithm and show that S3T has superior deletion performance compared to existing methods.
The rest of the paper is organized as follows: (a) We introduce the prior literature related to approximate and exact machine unlearning (Section 2), (b) We describe the problem setup for exact unlearning (Section 3.1), (c) We introduce the fine-tuning approach, S3T, and sequence selection algorithm under budget constraints (Section 3.23.3), (d) We theoretically analyze several properties of S3T (Section 3.4), and (e) We present experiments to evaluate the effectiveness of S3T’s fine-tuning approach, deletion performance and sequence selection algorithm (Section 4).
@@ 2 Background
Machine unlearning techniques [10] for deep learning can be broadly classified into two categories: approximate and exact unlearning. Approximate unlearning techniques focus on reducing the influence of a deleted instance from a model after it has been trained. Exact unlearning techniques provide unlearning guarantees by ensuring learning components trained on a deleted instance are not used during inference. In this section, we discuss each of these categories in detail.
@@
Approximate Machine Unlearning. These techniques focus on approximating the model parameters as if the deleted data instance was not there in the training set from the beginning [26]. These techniques typically quantify the influence of an instance [33] and perform gradient ascent for unlearning [22, 23, 45, 51, 27, 54, 40]. In contrast to these approaches, [25] stores the exact gradients encountered during training and uses them directly for gradient ascent. Another line of work [56, 57, 32, 11, 17, 48, 36, 41] focuses on unlearning in a batch setting, where they assume access to both a retention set and a forgetting set of data instances for approximate unlearning. While efficient in practice, auditing approximate unlearning techniques is challenging due to the stochastic nature of the optimization process [59, 63] and may have weak privacy guarantees in practice [28].
@@
Figure 1: Schematic diagram of the Sharded, Isolated, Sliced, and Aggregated training (SISA) [8] framework. An ensemble of models is individually trained on disjoint shards. (Left) Each shard is further divided into slices. (Right) Each model is sequentially trained on the slices and checkpoints are stored. After deletion, retraining resumes from the best available checkpoint.
@@
Exact Machine Unlearning. These techniques focus on developing a modular machine learning system, where individual components are trained using disjoint subsets of the data. Such a system offers the advantage that when a deletion request is received for an input instance, we only need to retrain the affected component rather than the entire model. The first instance of such a modular system: Sharded, Isolated, Sliced, and Aggregated training (SISA) [8]. SISA uses an ensemble of models each trained on a disjoint shard of the dataset as shown in Figure 1 (left). To further reduce retraining costs, each shard is divided into slices, and the models are incrementally trained on these slices, with their checkpoints stored sequentially (shown in Figure 1 (right)). If a deletion request is received for a data instance within 4th slice of a shard, we must retrieve the checkpoint from the 3rd slice and retrain using the remaining data within that shard. Several approaches focus on improving the components within SISA in application-specific settings like enhancing the dataset partitioning mechanism [3, 66], the retraining efficiency using light-weight adapters [35, 16], or extending the modular approach for vision tasks by using compartmentalized diffusion models [24]. However, these approaches are orthogonal to our work since they do not fundamentally modify the functioning of SISA and therefore can be easily incorporated within our proposed framework.
@@ 3.1 Problem Setting
We consider the general setting where the user fine-tunes a pre-trained model [14, 60] on private data using PEFT techniques. We assume that the deletion requests affect only the private fine-tuning data, not the pre-training data. Our framework, S3T, uses a setup where the dataset is divided into disjoint shards, with each shard further divided into slices. In S3T, we partition a dataset $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{m}\}$ into $m$ disjoint shards. Each shard is further divided into $L$ slices: $\mathcal{D}_{i}=\{S_{1},\ldots,S_{L}\}$. S3T trains a separate model for each shard and uses their aggregate decision.
Existing unlearning frameworks SISA [8] use a similar setup described above. In SISA, within each shard, the model is trained in multiple stages sequentially on the slices (training stages are Slice 1, Slice 1+2, and so on), and their checkpoints are stored. However, a key weakness of SISA is that individual models within the ensemble need to be retrained whenever a deletion request is executed. Even retraining on a single slice is expensive for large-scale models in production serving a huge customer base. A naive alternative would be to use the last known usable checkpoint and perform retraining after regular intervals. For example, in Figure 1 (right), if a deletion request arrives for a data instance in Slice 3, the model in production can be replaced with the checkpoint obtained after Slice 2. It is easy to see that the performance of the overall model will degrade with the number of deletion requests. In the extreme scenario where deletion requests target Slice 1 of all shards, the entire service goes down necessitating retraining from scratch.
We present an exact unlearning framework, S3T, to address these challenges. The core idea involves training several copies of the model with different slice sequences. When deletion requests occur, we utilize the model trained on a sequence that minimizes performance degradation. To further reduce the training cost, we leverage PEFT techniques and present a novel sequential slice-wise fine-tuning strategy in Section 3.2. Our training strategy allows us to use the same model by deactivating certain layers without the need to swap checkpoints in case of a deletion. In the following sections, we introduce the fine-tuning strategy and demonstrate how to select diverse data sequences for training.
@@ 3.2 Sequential Slice-wise Training
We describe our approach using the PEFT technique, LoRA [31], but our method is general can be easily extended to other PEFT techniques. LoRA introduces a small number of trainable low-rank ($r\ll d$) parameters, $(\bf X,\bf Y)$, while the pre-trained weights $\mathbf{\overline{W}}$ remains fixed as shown below:
@@
Our key idea involves training different LoRA layers using different slices of data. This approach allows us to selectively deactivate (zero out) the LoRA parameters ($\bf X$ or $\bf Y$) associated with a particular layer in the event of data deletion from that slice. In Figure 2 (left), we illustrate the training process in detail, where we follow a sequential top-to-bottom training approach. At stage 1, we train the final layer of the model (Layer 1 in the Figure) using data slice 1 while LoRA parameters from all other layers are switched off. In the next stage, we train Layer 2 using data slices 1 and 2, while keeping the LoRA parameters from the 1st layer frozen. This process continues for the rest of the layers. Note that our training process does not need to proceed in a layer-wise fashion, where each layer is trained on a different data slice. The user can even train $k$-LoRA layers per data slice resulting in a total of $kL$ layers being trained.
@@
The sequential slice-wise training process ensures that the LoRA parameter updates at the $i$-th layer are a function of the data instances within slices $\{1,\ldots,i\}$. Therefore, if a deletion request affects the $i$-th slice, the same model can still be used by deactivating the LoRA parameters corresponding to slices $\{i,\ldots,L\}$ (see details in Algorithm 3). For example, if a deletion request affects slice $S_{3}$ only the subsequent LoRA layers need to be deactivated to ensure exact unlearning, as shown in the first example of Figure 2 (right). This is because during training the parameters of layers 1 & 2 were not affected by instances in $S_{3}$. Now consider the scenario where deletion requests affect multiple slices. This is shown in the second example of Figure 2 (right), where slices $S_{1}$ and $S_{3}$ are affected. In this case, we observe that a model trained on the default ordering of the sequence $\{S_{1},\ldots,S_{L}\}$ is rendered useless when $S_{1}$ and $S_{3}$ are affected. This motivates us to train multiple models using different permutations of the slices. This would enhance the service time and system performance by selecting a model trained with the most effective ordering (e.g., the third ordering in Figure 2 (right) yields the best-performing model). However, training on all permutations of slices, $L!$, is expensive. In the following section 3.3, we present an algorithm to select a diverse set of permutations under resource constraints.
Design Choices. Here, we discuss the rationale behind the design choices for the proposed slice-wise training approach. First, we use top-to-bottom fine-tuning because we found the top layers were easier to train at the start and a bottom-up approach didn’t converge. Second, we train layers in a cumulative manner where the $i$-th layer is trained using slices $\{1,\ldots,i\}$. We found that this way of training helps in better convergence compared to the setup where we train every layer using a different slice. Third, we analyze the cost of slice-wise training and find it to be comparable to full training. Considering that training cost is $c=O(nl)$, where $n$ is the dataset size and $l$ is the number of trainable layers (empirical evidence in Appendix C.2). For slice-wise training, we observe that $c=\sum_{i=1}^{k}(\frac{in}{k})(\frac{l}{k})=O\left(\frac{nl}{2}\right)$, which is of the same order as full training.
@@
Figure 3: Illustration of the slice sequence selection problem under a budget constraint. (Left) An example of a permutation tree with $L=3$ and a diverse set of sequences for budget $B=3$ is shown in green. (Center) We show the intuition behind the BMS algorithm, where a bipartite graph is formed between successive layers, and we find a perfect matching between them. (Right) Illustration of the conditional sampling, where each slice is selected sequentially based on the deletion probabilities.

### v2

@@ Abstract
Machine unlearning is the process of efficiently removing the influence of a training data instance from a trained machine learning model without retraining it from scratch. A popular subclass of unlearning approaches is exact machine unlearning, which focuses on techniques that explicitly guarantee the removal of the influence of a data instance from a model. Exact unlearning approaches use a machine learning model in which individual components are trained on disjoint subsets of the data. During deletion, exact unlearning approaches only retrain the affected components rather than the entire model. While existing approaches reduce retraining costs, it can still be expensive for an organization to retrain a model component as it requires halting a system in production, which leads to service failure and adversely impacts customers. To address these challenges, we introduce an exact unlearning framework – Sequence-aware Sharded Sliced Training (S3T), which is designed to enhance the deletion capabilities of an exact unlearning system while minimizing the impact on model’s performance. At the core of S3T, we utilize a lightweight parameter-efficient fine-tuning approach that enables parameter isolation by sequentially training layers with disjoint data slices. This enables efficient unlearning by simply deactivating the layers affected by data deletion. Furthermore, to reduce the retraining cost and improve model performance, we train the model on multiple data sequences, which allows S3T to handle an increased number of deletion requests. Both theoretically and empirically, we demonstrate that S3T attains superior deletion capabilities and enhanced performance compared to baselines across a wide range of settings.
@@ 1 Introduction
In recent years, the growing success of machine learning (ML) has led to its widespread deployment across a range of applications (Achiam et al., 2023; Team et al., 2023; Qayyum et al., 2020; Surden, 2021). Once a machine learning model has been trained, it is often necessary to unlearn specific training data instances for various reasons, like complying with user data deletion requests (Mantelero, 2013; European Parliament & Council of the European Union, ; Shastri et al., 2019; Achille et al., 2024), removing stale or corrupt data (Biggio et al., 2012; Steinhardt et al., 2017), etc. Retraining an ML model entirely from scratch with each deletion request is expensive, especially for modern large-scale models (Brown et al., 2020; Achiam et al., 2023; Team et al., 2023). Machine unlearning (Nguyen et al., 2022; Xu et al., 2023) techniques focus on efficiently unlearning the influence of a data instance from a trained machine learning model.
@@
Machine unlearning techniques are classified into two broad categories: approximate and exact unlearning (Xu et al., 2024). Approximate unlearning techniques (Guo et al., 2020; Liu et al., 2024a) modify the parameters of a trained model to reduce the influence of the deleted data instance. While cost-effective in practice, approximate unlearning cannot guarantee the complete removal of an instance’s influence and it may still retain non-zero influence on the model. Moreover, auditing approximate unlearning is challenging due to the stochastic nature of ML optimization (Thudi et al., 2022). An alternative to this approach is exact unlearning (Cao & Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2023), which can guarantee the removal of a data instance’s influence from a trained model. Exact unlearning techniques use a modular system, where different components within the system are trained on disjoint data subsets (Section 2). When a deletion request occurs, only the affected component needs to be retrained. However, in real-world settings, halting a production system to even retrain a single component can result in service failure and significant expenses. The alternative is to function without the affected component, which may result in reduced performance, ultimately impacting end users. To address these challenges, we introduce Sequence-aware Sharded Sliced Training (S3T), which can exactly unlearn private fine-tuning data efficiently while minimizing the impact on performance.
@@
The key idea behind our S3T framework is to perform additional offline training before deploying the initial model to reduce retraining costs. At the core of S3T, we leverage a novel lightweight fine-tuning approach that allows parameter isolation by sequentially training model layers using disjoint data slices (Section 3.2). Due to this parameter isolation, it is possible to efficiently perform exact unlearning by deactivating the layers associated with the deleted instance, rather than discarding the entire checkpoint. We efficiently train multiple models using different sequences of the same data slices depending on a training budget. The training budget decides the number of models we can train. We show that increasing the training budget before deployment can significantly reduce retraining costs and improve the model’s performance. We also observe that it is important to train the model using diverse sequences and provide several approaches for selecting diverse sequences (Section 3.3). Furthermore, we formalize the notion of deletion rate for exact unlearning systems and theoretically show that S3T achieves provably better deletion guarantees than existing approaches (Section 3.4). We conduct extensive empirical evaluations to evaluate the effectiveness of S3T using Transformers with parameter counts ranging from 86M to 13B on a range of tasks. Finally, in Section 4, we empirically validate the sequence selection algorithm and show that S3T has superior deletion performance compared to existing methods in real-world settings.
@@ 2 Background
Machine unlearning techniques for deep learning is broadly classified into two categories: approximate and exact unlearning (Xu et al., 2024). Approximate unlearning techniques focus on reducing the influence of a deleted instance from a model after it has been trained. Exact unlearning techniques provide unlearning guarantees by ensuring model components trained on a deleted instance are not used during inference. In this section, we discuss each of these categories in detail.
@@
Approximate Machine Unlearning. These techniques focus on approximating the model parameters as if the deleted data instance was not there in the training set from the beginning (Guo et al., 2020). These techniques typically quantify the influence of an instance (Koh & Liang, 2017) and perform gradient ascent for unlearning (Golatkar et al., 2020a; b; Neel et al., 2021; Sekhari et al., 2021; Gupta et al., 2021; Suriyakumar & Wilson, 2022; Liu et al., 2024a). In contrast to these approaches, (Graves et al., 2021) stores the exact gradients encountered during training and uses them directly for gradient ascent. Another line of work (Tarun et al., 2023a; b; Jia et al., 2023; Chen & Yang, 2023; Eldan & Russinovich, 2023; Patil et al., 2023; Kurmanji et al., 2024; Liu et al., 2024b; Park et al., 2024; Zhang et al., 2024) focuses on unlearning in a batch setting, where they assume access to both a retention set and a forget set of data instances for approximate unlearning. While efficient in practice, auditing approximate unlearning techniques is challenging due to the stochastic nature of the optimization process (Thudi et al., 2022; Wang et al., 2024) and may have weak privacy guarantees in practice (Hayes et al., 2024; Łucki et al., 2024).
@@
Figure 1: Schematic diagram of the Sharded, Isolated, Sliced, and Aggregated training (SISA) (Bourtoule et al., 2021) framework. An ensemble of models is individually trained on disjoint shards. (Left) Each shard is further divided into slices. (Right) Each model is sequentially trained on the slices and checkpoints are stored. After deletion, retraining resumes from the best available checkpoint.
@@
Exact Machine Unlearning. These techniques focus on developing a modular machine learning system, where individual components are trained using disjoint subsets of the data. Such a system offers the advantage that when a deletion request is received for an input instance, we only need to retrain the affected component rather than the entire model. However, these systems require modifying the original training process of the model. The seminal work for such a modular unlearning system is Sharded, Isolated, Sliced, and Aggregated training (SISA) (Bourtoule et al., 2021). SISA uses an ensemble of models each trained on a disjoint shard of the dataset as shown in Figure 1 (left). To further reduce retraining costs, each shard is divided into slices, and the models are incrementally trained on these slices, with their checkpoints stored sequentially (shown in Figure 1 (right)). If a deletion request affects a data instance within the 4th slice of a shard, we must retrieve the checkpoint after the 3rd slice and retrain using the remaining data within that shard. Several approaches focus on improving the components within SISA in application-specific settings like enhancing the dataset partitioning mechanism (Aldaghri et al., 2021; Yan et al., 2022), the retraining efficiency using light-weight adapters (Kumar et al., 2023; Dukler et al., 2023), or extending the modular approach for vision tasks by using compartmentalized diffusion models (Golatkar et al., 2023). However, these approaches are orthogonal to our work since they do not fundamentally modify the functioning of SISA. These modifications can be easily incorporated within our proposed framework.
@@ 3.1 Problem Setting
We consider the general setting where the user fine-tunes a pre-trained model like BERT (Devlin et al., 2019) or Llama (Touvron et al., 2023) on private data using PEFT techniques. We assume that the deletion requests affect only the private fine-tuning data, not the pre-training data. In S3T, we partition a dataset $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{m}\}$ into $m$ disjoint shards. Each shard is further divided into $L$ slices: $\mathcal{D}_{i}=\{S_{1},\ldots,S_{L}\}$. S3T trains a separate model per shard and uses their aggregate decision.
Existing unlearning frameworks SISA (Bourtoule et al., 2021) use a similar setup described above. In SISA, within each shard, the model is trained in multiple stages sequentially on the slices (training stages are Slice 1, Slice 1+2, and so on), and their checkpoints are stored. However, a key weakness of SISA is that if deletion requests affect Slice 1 of all shards, then the entire service goes down necessitating retraining from scratch. Another drawback of SISA is that individual models within the ensemble need to be retrained whenever a deletion request is executed. Retraining even on a single slice is expensive for large-scale models in production serving a huge customer base. A naive alternative would be to use the last known usable checkpoint and perform retraining after regular intervals. For example, in Figure 1 (right), if a deletion request arrives for a data instance in Slice 4, the model in production can be replaced with the checkpoint obtained after Slice 3. It is easy to see that the performance of the overall model will degrade with the number of deletion requests.
We present an exact unlearning framework, S3T, to address these challenges. The core idea involves training several copies of each model (within the ensemble) using different slice sequences. When deletion requests occur, we utilize the model that minimizes performance degradation. To further reduce the training cost, we leverage PEFT techniques and present a novel sequential slice-wise fine-tuning strategy in Section 3.2. Our training strategy allows us to use the same model by deactivating certain layers without the need to swap checkpoints in case of a deletion. In the following sections, we introduce the fine-tuning strategy and sequence selection process.
@@ 3.2 Sequential Slice-wise Training
We describe S3T using the PEFT technique, LoRA (Hu et al., 2021), but our method is general can be easily extended to other PEFT techniques. LoRA introduces a small number of trainable low-rank ($r\ll d$) parameters, $(\bf X,\bf Y)$, while the pre-trained weights $\mathbf{\overline{W}}$ remains fixed as shown below:
@@
Our key idea involves training different LoRA layers using different data slices. This approach allows us to selectively deactivate (zero out) the LoRA parameters ($\bf X$ or $\bf Y$) associated with a particular layer in the event of data deletion from that slice. In Figure 2 (left), we illustrate the training process in detail, where we follow a sequential top-to-bottom training approach. At stage 1, we train the final model layer (Layer 1 in the figure) using slice 1 while LoRA parameters from all other layers are switched off. In the next stage, we train second last layer (Layer 2) using slices 1 & 2, while keeping the LoRA parameters from the Layer 1 frozen. This process continues for the rest of the layers. Note that the training does not need to proceed in a single layer-wise fashion, we can even train multiple LoRA layers per slice. We discuss more details about the design choice in Appendix C.2.
@@
The sequential slice-wise training process ensures that the LoRA parameter updates at the $i$-th layer are a function of the data instances within slices $\{1,\ldots,i\}$. Therefore, if a deletion request affects the $i$-th slice, the same model can still be used by deactivating the LoRA parameters corresponding to slices $\{i,\ldots,L\}$ (see details in Algorithm 4). For example, if a deletion request affects slice $S_{3}$ only the subsequent LoRA layers need to be deactivated to ensure exact unlearning, as shown in the first example of Figure 2 (right). This is because during training the parameters of layers 1 & 2 were not affected by instances in $S_{3}$. During this deletion process, we use the same model checkpoint and switch off LoRA layers, resulting in an $L$-time reduction in storage cost compared to SISA.
Now we consider the scenario where deletion requests affect multiple slices. This is shown in the 2nd sequence of Figure 2 (right), where slices $S_{1}$ and $S_{3}$ are affected. In this case, we observe that a model trained on the default ordering of the sequence $\{S_{1},\ldots,S_{5}\}$ is rendered useless when $S_{1}$ and $S_{3}$ are affected. This motivates us to train multiple models using different permutations of the slices. This would enhance the service time and system performance by selecting a model trained with the most effective ordering (e.g., the 3rd sequence in Figure 2 (right) yields the best-performing model). However, training on all $L!$ slice permutations is prohibitively expensive. In the following section 3.3, we present strategies to select a diverse set of permutations under budget constraints. Using these selection approaches, in Section 3.4 we theoretically show that we do not need more than $L$ sequences to achieve the optimal deletion performance.
@@
Figure 3: Illustration of the slice sequence selection problem with uniform deletion prior under a budget constraint, $B$. (Left) An example of a permutation tree with $L=3$ and a diverse set of sequences for budget $B=3$ is shown in green. (Center) We show the functioning of the cyclic rotation algorithm, where we generate cyclic permutations of the original sequence. (Right) We iteratively extend the algorithm when budget $B>L$ by generating cyclic rotations of the subsequences.